Simplify the following, using the rules of exponents

i) x9. x8

ii) x5 . x-9

iii) x-11. x

iv) x6 . x0.15

v)𝑥5/𝑥−3

vi)𝑥−4/𝑥−7

vii)𝑥−3/𝑥−4

viii)𝑥9/√𝑥

ix) (𝑥−4)−5

x) (𝑥−0.4)−2

xi)1/𝑥−1.5. 1/𝑦−2.5

xii)𝑥−3.5/𝑦−2.2



1
Expert's answer
2021-09-09T11:04:58-0400

Following below are some common rules of exponent:

1. Product rule"= a^m \u00d7an^ = a^{m+n}"

2. Division rule "= a^m\u00f7a^n = a^{m\u2212n}"

3. Zero exponents "= a^0 = 1"

4. Negative exponents "= a^{−m} = \frac{1}{a^m}"


"i) x^9. x^8\\
= X^{9+8}\\ 
= X^{17}"


"ii) x^5. x^{−9}\\
= X^{5+(−9)}\\
= X^{5−9}\\
= X^{−4}"

"iii) x^{−11}. x \\
= X^{−11+1}\\
= X^{−10}"